2. What was the expected volume of hydrogen produced in the reaction in part 2? Compare your result with the value you measured in the lab and calculate the percent error using the formula below. l experimental volume calculated volume calculated volume e % error = x 100

Data related to the question:



If any other information is required, let me know. Thanks!
Answer #1

Well done! You have performed almost all the calculations required to solve the problem. The rest can be solved as

At STP (Standard Temperature and Pressure, i.e. T = 273.15 K and P = 1 atm), the molar volume of any gas (Here, hydrogen gas) is given by 22.4 L, i.e. 22.4 L/mol. (Expected value)

The calculated value for the molar volume of hydrogen gas = (24.14 + 24.12)/2, i.e. 24.13 L, i.e. 24.13 L/mol.

The difference between the calculated and expected values = 24.13 - 22.4, i.e. 1.73 L.

Therefore, the % error = 1.73 * 100 / 22.4, i.e. 7.72.